# uncompress

### v1

@@ 5.4 Limitations
Alternate natural language processing (NLP) decisions would change our results. In all data cleaning procedures, we maintain as much of the original data as feasible, and we run parallel analyses on more- or less-cleaned data to inform our choices, which impact the distributions of token, part of speech, and interjection frequencies. We run our analyses with and without punctuation, ultimately relying of analyses with punctuation; punctuation can convey content from prosody [wolf2023prosody_text_redundancy], and including it with the rest of text can more completely express Zipf’s law [stanisz2024nlp_complexity]. When important, we note throughout our work the substantive differences we find between these two instances of the data (without versus with punctuation). Another possible limitation arises from cleaning the mediums to be more comparable, such as the removal of grammatical artifacts from fictional dialogues under the assumption that these would not be acted or spoken. These NLP choices could exclude the unique signal found in movie conversations. A future project could investigate whether there are meaningful effects of NLP and medium on the fit of Heaps’ law by comparing additional data cleaning pipelines applied to the same corpora. Likewise, our choice of conversation segmentation algorithm shapes the data. Research could compare conversation segmentation algorithms to see how they might change the shape of conversations.
@@ 5.5 Future work
Conversation research can gain further meaning from text. First, conversation topics would be fruitful in identifying meaning. Regression coefficients reflecting Heaps’ law scaling and cosine similarity are simple ways to start exploring how to identify abrupt and meaningful topic shifts and callbacks across conversation windows. These minimalist topic markers could provide context and be compared or used in conjunction with models such as BERTopic. Second, interjections offer rich signal that could be furhter disambiguated by type. A common type distinction is interjection words versus interjectional phrases, often thought of as parts of speech versus syntax. Interjections whose primary role is feedback can exhibit uncommon letter sequences and sounds (e.g., pfft, tut-tut[ameka1992interjections]; paired with the syntactic construction of interjectional phrases, interjection forms have tagging and token implications and may show different scaling between types when correctly recognized. Interlocutor feedback takes various forms—including the lexical form on which we focus—and fulfills many roles, such as that of continuers, expressing a reaction, and acknowledgment. Consistent with this purpose, we find interjections’ appearance cadence in naturalistic conversation to be more regular than in scripted conversation, and we find more use of typical filler words in naturalistic speech. Generic feedback, such as interjection words, likely occur during conversational set-up, whereas specific feedback (such as interjectional phrases) is more closely connected to the local context [boudin2024multimodal_feedback_in_convo]. Future research could pair the timing of interjections with their word-forms to qualify when the purpose of their usage changes (such as simply showing interest, good manner or showing common ground).
@@ Appendix A Appendices
We refrain from mentioning specific content as researchers must register a request to access teh data. Further investigation of two topics found in the text showed that even of those 2 highly-mentioned topics, the 2 participants making those mentions the most only made them 20 and 26 times in 6 and 4 conversations respectively. The two topics appeared each in less than 1% of utterances in CANDOR. At the maximum, an outsized impact of speakers and a practice effect could look like 3 speakers mentioning the same topics in 8 conversations at a higher rate.
@@
We offer two methodological notes. First consider documents lengths: Movies (I) distributions are artificially shorter because the text lengths are much shorter. Document length also affects seeing extremely high arrival times of unique types because longer texts have more time for novel types to occur. Second, calculating burstiness requires more than 2 items of the same kind because 2 items provides only 1 interval that results in mean and standard deviation of 0 and a burstiness value of -1, stacking extremely rare words in this region. We filter these rare word occurrences when calculating at the corpora level (on words), but this filtering inherently does not happen for POS because there are 5 categories that are filled every time and do not have rare types. Thus, the POS and corpora-level B and M averages are slightly reflective of this difference while the corpora and literary cluster may be easier to compare.
@@
Figure A5: Interarrival times by corpus and part of speech in log10 scale. The left column is total tokens while the right column is unique tokens.
@@
Figure A6: Burstiness and memory by corpora and parts of speech. Top: Burstiness and memory by corpora and parts of speech. Left is the full metric range from -1 to 1, and the right plot zooms into the data range. Asterisked labels are those corpora shuffled. Bottom: Burstiness and memory by corpora and part of speech with values reported as ‘B / M’. Please refer to the methodological note in this section about comparing corpora averages and POS.
@@
We recode raw POS tags using the following schema [chacoma_tags_2020]: adverb $\rightarrow$ verb, pronoun $\rightarrow$ noun, proper noun $\rightarrow$ noun, adposition $\rightarrow$ function words, coordinating conjunction $\rightarrow$ function words, subordinating conjunction $\rightarrow$ function words, interjections remained interjections, and all remaining tags $\rightarrow$ other.
@@
We find some erroneously recorded start/stop times that result in implausible pauses (approximately 25 minutes). These could be identified because the recording start/time time jumps for just that utterance or the pause does not reflect the start/stop time difference. At least 6 conversations were affected at a minimum.
@@
Note that the range for pauses (in seconds) is negative to positive, reflecting overlap in speech (interruptions) or otherwise gaps between speakers’ turns. Some utterances have a list of pauses which seems to reflect the timings around a listener’s parallel speech. These are not representative of pauses’ overlaps (interruptions) or gaps between turns, so we remove these pauses for reporting. In fact, with these pauses included, the average is negative ($\mu=-0.22,~\text{median}=1.18,~\sigma=15.24$) but positive with parallel speech pauses removed ($\mu=1.14,~\text{median}=0.61,~\sigma=8.04$). We find that these statistics are sensitive to the technical problems’ outliers, hence including the median. Prior CANDOR work [reece2023candor] using the Backbiter model showed the average pause interval was 0.2 seconds ($\text{median}=0.38$ seconds) with gaps largely under 1 second. The difference between NaturalTurn and Backbiter is that NaturalTurn keeps a speaker’s turn together while preserving concurrent parallel speech, but the parallel speech ‘interruptions’ do not intersperse the speaker’s coherent turn; thus, the length of utterances and pauses may differ from those reported under Backbiter. Given that online connection issues seem present in the turn segmenting, it tended to stretch the mean and standard deviation, so we suggest more reliance on the median and methods robust to outliers.
@@
Next are conversation-level distributions for backchannels, secondary speech, secondary + backchannel occurrences, and pauses. We normalize these occurrences by number of utterances in the conversation. We quantify the spread of spread of these parallel speech features, which could inform future work related to parallel speech; parallel speech includes language that mirror’s a storyteller’s emotion or reinforces key moments [cooney2024naturalturn]—both of which can be pro-social in nature.
@@
When pause scores in a conversation are normalized by the amount of utterances, the distribution becomes normal (counter to its skew in both raw amount of pauses and unnormalized pause scores per conversation). This normal appearance may support some adherence to pauses that feel more tenable or desirable in a conversation. Further work should connect scores such as these to quality outcomes. It would also be interesting to asses response gaps in non-naturalistic conversation mediums such as online to see if the same perception exists regarding time-to-reply gaps.

### v2

@@ 5.4 Limitations
Alternate NLP decisions would change our results. In all data cleaning procedures, we maintain as much of the original data as feasible, and we run parallel analyses on more- or less-cleaned data to inform our choices, which impact the distributions of token, part of speech, and interjection frequencies. We run our analyses with and without punctuation, ultimately relying on analyses with punctuation; punctuation can convey content from prosody [wolf2023prosody_text_redundancy], and including it with the rest of text can more completely express Zipf’s law [stanisz2024nlp_complexity]. When important, we note throughout our work the substantive differences we find between these two instances of the data (without versus with punctuation). Another possible limitation arises from cleaning the mediums to be more comparable, such as the removal of grammatical artifacts from fictional dialogues under the assumption that these would not be acted or spoken. These NLP choices could exclude the unique signal found in movie conversations. A future project could investigate whether there are meaningful effects of NLP and medium on the fit of Heaps’ law by comparing additional data cleaning pipelines applied to the same corpora. Likewise, our choice of conversation segmentation algorithm shapes the data. Research could compare conversation segmentation algorithms to see how they might change the shape of conversations.
@@ 5.5 Future work
Conversation research can gain further meaning from text. First, conversation topics would be fruitful in identifying meaning. Regression coefficients reflecting Heaps’ law scaling and cosine similarity are simple ways to start exploring how to identify abrupt and meaningful topic shifts and callbacks across conversation windows. These minimalist topic markers could provide context and be compared or used in conjunction with models such as BERTopic. Second, interjections offer rich signal that could be further disambiguated by type. A common type distinction is interjection words versus interjectional phrases, often thought of as parts of speech versus syntax. Interjections whose primary role is feedback can exhibit uncommon letter sequences and sounds (e.g., pfft, tut-tut[ameka1992interjections]; paired with the syntactic construction of interjectional phrases, interjection forms have tagging and token implications and may show different scaling between types when correctly recognized. Interlocutor feedback takes various forms—including the lexical form on which we focus—and fulfills many roles, such as that of continuers, expressing a reaction, and acknowledgment. Consistent with this purpose, we find interjections’ appearance cadence in naturalistic conversation to be more regular than in scripted conversation, and we find more use of typical filler words in naturalistic speech. Generic feedback, such as interjection words, likely occurs during conversational set-up, whereas specific feedback (such as interjectional phrases) is more closely connected to the local context [boudin2024multimodal_feedback_in_convo]. Future research could pair the timing of interjections with their word-forms to qualify when the purpose of their usage changes (such as simply showing interest, good manner or showing common ground).
@@ Appendix A Appendices
We refrain from mentioning specific content as researchers must register a request to access the data. Further investigation of two topics found in the text showed that even of those 2 highly-mentioned topics, the 2 participants making those mentions the most only made them 20 and 26 times in 6 and 4 conversations respectively. The two topics appeared each in less than 1% of utterances in CANDOR. At the maximum, an outsized impact of speakers and a practice effect could look like 3 speakers mentioning the same topics in 8 conversations at a higher rate.
@@
We offer two methodological notes. First consider documents lengths: Movies (I) distributions are artificially shorter because the text lengths are much shorter. Document length also affects seeing extremely high arrival times of unique types because longer texts have more time for novel types to occur. Second, calculating burstiness requires more than 2 items of the same kind because 2 items provide only 1 interval that results in mean and standard deviation of 0 and a burstiness value of -1, stacking extremely rare words in this region. We filter these rare word occurrences when calculating at the corpora level (on words), but this filtering inherently does not happen for POS because there are 5 categories that are filled every time and do not have rare types. Thus, the POS and corpora-level B and M averages are slightly reflective of this difference while the corpora and literary cluster may be easier to compare.
@@
Figure A5: Interarrival times by corpus and part of speech in log10 scale. The left column represents total tokens while the right column represents unique tokens.
@@
Figure A6: Burstiness and memory by corpora and parts of speech. Top: Burstiness and memory by corpora and parts of speech. Left is the full metric range from -1 to 1, and the right plot zooms into the data range. Asterisked labels are those corpora shuffled. Bottom: Burstiness and memory by corpora and part of speech with values reported as $\text{B}/\text{M}$. Please refer to the methodological note in this section about comparing corpora averages and POS.
@@
We recode raw POS tags using the following schema [chacoma_tags_2020]: adverb $\rightarrow$ verb, pronoun $\rightarrow$ noun, proper noun $\rightarrow$ noun, adposition $\rightarrow$ function words, coordinating conjunction $\rightarrow$ function words, subordinating conjunction $\rightarrow$ function words, interjections remain interjections, and all remaining tags $\rightarrow$ other.
@@
We find some erroneously recorded start/stop times that result in implausible pauses (approximately 25 minutes). These could be identified because the recording start/stop time jumps for just that utterance or the pause does not reflect the start/stop time difference. At least 6 conversations were affected at a minimum.
@@
Note that the range for pauses (in seconds) is negative to positive, reflecting overlap in speech (interruptions) or otherwise gaps between speakers’ turns. Some utterances have a list of pauses which seems to reflect the timings around a listener’s parallel speech. These are not representative of pauses’ overlaps (interruptions) or gaps between turns, so we remove these pauses for reporting. In fact, with these pauses included, the average is negative ($\mu=-0.22,~\text{median}=1.18,~\sigma=15.24$) but positive with parallel speech pauses removed ($\mu=1.14,~\text{median}=0.61,~\sigma=8.04$). We find that these statistics are sensitive to the technical problems’ outliers, hence including the median. Prior CANDOR work [reece2023candor] using the Backbiter model showed the average pause interval was 0.2 seconds ($\text{median}=0.38$ seconds) with gaps largely under 1 second. The difference between NaturalTurn and Backbiter is that NaturalTurn keeps a speaker’s turn together while preserving concurrent parallel speech, but the parallel speech ‘interruptions’ do not intersperse the speaker’s coherent turn; thus, the length of utterances and pauses may differ from those reported under Backbiter. Given that online connection issues seem present in the turn segmenting and tended to stretch the mean and standard deviation, we suggest more reliance on the median and methods robust to outliers.
@@
Next are conversation-level distributions for backchannels, secondary speech, secondary + backchannel occurrences, and pauses. We normalize these occurrences by number of utterances in the conversation. We quantify the spread of these parallel speech features, which could inform future work related to parallel speech; parallel speech includes language that mirrors a storyteller’s emotion or reinforces key moments [cooney2024naturalturn]—both of which can be pro-social in nature.
@@
When pause scores in a conversation are normalized by the amount of utterances, the distribution becomes normal (counter to its skew in both raw amount of pauses and unnormalized pause scores per conversation). This normal appearance may support some adherence to pauses that feel more tenable or desirable in a conversation. Further work should connect scores such as these to quality outcomes. It would also be interesting to assess response gaps in non-naturalistic conversation mediums such as online to see if the same perception exists regarding time-to-reply gaps.